write the equation of the line that passes through point (6,-7) with a slope of -2

Answer by checkley71(8403)

Y=mX+b
-7=-2*6+b
-7=-12+b
b=-7+12
b=5 ANSWER FOR THE Y INTERCEPT.
THUS THE EQUATION IS:
Y=-2X+5 ANSWER.